Old Time Savings Bank pays 4% interest on its savings accounts. If you deposit $1,000 in the bank and leave it there: (Do not round intermediate calculations. Round your answers to 2 decimal places.)

a. How much interest will you earn in the second year?

b. How much interest will you earn in the 10th year?
Answer #1

Amount=P*(1+r)^Time period

P is the principal amount and r is the interest rate

Part a:
Amount=P*(1+r)^Time period
When p=$1000, r=4% and time period is 2 years
Amount in second year=1000*(1+4%)^2=1081.6

Interest amount=1081.6-1000=$81.60

Part b:
Amount=P*(1+r)^Time period
Amount in tenth year=1000*(1+4%)^10=1480.24
Interest amount=1480.24-1000=$480.24